One tablespoon of peanut butter has a mass of 17 g
Answer #1

The heat released by the 17 grams of PB is Q = 17 x 110 x (25.2--22)

                                                                  Q = 352 kj

If we convert to joules Q = 352 kJ = 352,000 J

Calculate the released heat per gram = 352000 /17 = 20705 J/ g

4.18- J of thermal energy is equivalent to 1 calorie, divide the J/g by 4.184 to get calories/g

20705 / 4.184 = 4948.8 C

1000 calories = 1 food Calorie

Convert to food calorie = 4948.8 /1000 = 4.948.8 Cal/g